A 273 mL aqueous solution of NaBr has an osmotic pressure of 1.98 x 10-2 atm at a temperature of 45°C. Calculate the molarity of the solution.
1
Expert's answer
2020-04-23T13:08:47-0400

According to the osmotic pressure equation:

π = iMRT

where pi - osmotic pressure, i - van't Hoff's factor (number of ions formed in the solution), M - molarity, R gas constant, T - temperature.

From here:

M = π / (iRT)

As R = 0.08206 L atm / mol K, T = 45°C = 318.15 K, and i = 2 (NaBr dissociates completely producing two ions):

M = 1.98 × 10-2 atm / (2 × 0.08206 L atm / mol K × 318.15 K) = 3.8 × 10-4 mol/L


Answer: 3.8 × 10-4 mol/L